OpenStudy (anonymous):

The excess of an angle over its supplement is 20 degrees. Find the angle.

OpenStudy (anonymous):

100 degrees...

OpenStudy (anonymous):

you arent using the right equation..

OpenStudy (anonymous):

let n= the angle 180-n+ the supp that what your supposed to use..

OpenStudy (anonymous):

180-n= the supp *

OpenStudy (anonymous):

This is how I'm reading it. 100 and 80 are supplementary to each other or as you put it 180 - 100 = 80 and 180 - 80 =100 100 has an excess of 20 degrees over 80 so that's why I said 100 degrees.